# Analysis of Hessian Scaling for Local and Global Costs in Variational Quantum Algorithms

### v1

@@ 3.2 Main result
Assumption 3.2 concerns second-moment concentration of the objective value at random initialization. The following analysis shows that such concentration is stable under constant-size finite-shift operators; consequently, Hessian entries computed via parameter-shift rules inherit the same exponential rate, up to constants determined by the shift coefficients (Lemma 3.1). This can be justified in two ways. (i) If the induced state ensemble forms an approximate projective $2$-design, then concentration follows; Appendix B.2 proves $\operatorname{Var}_{\rho}[C(\bm{\theta})]\leq 1/(2^{n}+1)$ for $\|O\|\leq 1$ (cf. Proposition B.3). (ii) For modest-depth ansatze without such guarantees, (15) is directly testable via Monte Carlo estimates of $\operatorname{Var}_{\rho}[C(\bm{\theta})]$; by shift-invariance of $\rho$, the same estimate applies uniformly over the finite shift sets used in parameter-shift rules. We numerically verify this assumption in § 4.2 and give more discussion in Appendix B.2.
The next lemma records shift-uniformity and finite-shift stability of second-moment concentration, which will be used for global objectives.
@@
Let $\rho$ be shift-invariant on the torus, i.e., $\theta\sim\rho\Rightarrow\theta\oplus s\sim\rho$ for all shifts $s\in\mathbb{R}^{M}$. Fix a finite shift set $S\subset\mathbb{R}^{M}$ and coefficients $\{a_{s}\}_{s\in S}\subset\mathbb{R}$, and define $X_{s}:=C(\theta\oplus s)$. If $\operatorname{Var}_{\rho}[C(\theta)]\leq v_{n}$, then the same bound holds uniformly over shifts,
$\operatorname{Var}_{\rho}[X_{s}]=\operatorname{Var}_{\rho}[C(\theta\oplus s)]=\operatorname{Var}_{\rho}[C(\theta)]\leq v_{n}\qquad\text{for all }s\in S,$
@@
$\operatorname{Var}_{\rho}\Big(\sum_{s\in S}a_{s}C(\theta\oplus s)\Big)\leq\Big(\sum_{s\in S}|a_{s}|\Big)^{2}v_{n}.$
Proof 3.2.
Shift-invariance implies $\theta$ and $\theta\oplus s$ have the same distribution under $\rho$, hence $\operatorname{Var}_{\rho}[C(\theta\oplus s)]=\operatorname{Var}_{\rho}[C(\theta)]$ for all $s\in S$. For the linear combination, by bilinearity,
$\operatorname{Var}\Big(\sum_{s\in S}a_{s}X_{s}\Big)=\sum_{s,s^{\prime}\in S}a_{s}a_{s^{\prime}}\operatorname{Cov}(X_{s},X_{s^{\prime}}).$
@@
$\operatorname{Var}\Big(\sum_{s\in S}a_{s}X_{s}\Big)\leq v_{n}\sum_{s,s^{\prime}\in S}|a_{s}||a_{s^{\prime}}|=v_{n}\Big(\sum_{s\in S}|a_{s}|\Big)^{2}.$
@@
(16) $V_{G}(m):=\max_{v\in V(G)}|B_{G}(v,m)|.$
For bounded-degree graphs, $V_{G}(m)$ is finite for each fixed $m$; for $D$-dimensional lattices one has $V_{G}(m)=\mathcal{O}(m^{D})$.
@@
Proof 3.4 (Proof sketch).
@@
For global objectives, Lemma 3.1 transfers the assumed second-moment concentration of $C(\bm{\theta})$ uniformly over the finite shifts and preserves the exponential rate under any $O(1)$-shift finite-difference operator, yielding (17).
For local objectives, bounded lightcone growth implies that local contributions with sufficiently separated supports are independent; a dependency-graph argument then gives $\operatorname{Var}_{\rho}(C(\bm{\theta}\oplus s))=\mathcal{O}(V_{G}(k+2rL)/n)$ uniformly in $s$. Plugging this bound into the covariance representation yields (18). Details are deferred to Appendix A.3.
@@
The estimate (18) is an upper bound obtained via a generic dependency-graph argument and is not expected to be exponent-tight in typical instances. The key of Theorem 3.3 is the polynomial–versus–exponential resolvability dichotomy and its dependence on lightcone growth, rather than the optimal polynomial exponent.
Theorem 3.3 establishes a sharp dichotomy at initialization: termwise local averaging with sub-extensive lightcone growth leads to polynomially decaying variance, whereas global objectives satisfying Assumption 3.2 exhibit exponential decay. These results explicitly characterize the statistical resolvability of individual Hessian entries.
3.3 From entrywise variance to matrix-scale bounds and sample complexity
While the entrywise bounds above quantify the typical fluctuation scale, the practical utility of curvature information depends on the measurement budget required to separate these fluctuations from statistical noise, and on how they aggregate into matrix-level properties. Throughout this subsection, $\mathbb{E}_{\rho}[\cdot]$ and $\operatorname{Var}_{\rho}(\cdot)$ denote expectation and variance with respect to initialization $\bm{\theta}\sim\rho$. When finite-shot estimation is involved, we write $\mathbb{E}_{\mathrm{sh}}[\cdot\mid\bm{\theta}]$ and $\operatorname{Var}_{\mathrm{sh}}(\cdot\mid\bm{\theta})$ for expectation and variance over measurement noise (shots), conditional on a fixed parameter value $\bm{\theta}$.
We first record the standard implication of entrywise second-moment control for matrix norms. The proof is included for completeness in Appendix B.3.
Lemma 3.7.
For a random matrix $H\in\mathbb{R}^{M\times M}$ with finite second moments,
$\mathbb{E}\|H\|_{F}^{2}=\sum_{j,k=1}^{M}\Big(\operatorname{Var}(H_{jk})+(\mathbb{E}[H_{jk}])^{2}\Big),\quad\mathbb{E}\|H\|_{2}^{2}\leq\mathbb{E}\|H\|_{F}^{2},\quad\mathbb{E}\|H\|_{2}\leq\sqrt{\mathbb{E}\|H\|_{F}^{2}}.$
Consequently, if $\operatorname{Var}(H_{jk})\leq v_{n}$ and $|\mathbb{E}H_{jk}|\leq\mu_{n}$ for all $j,k$, then
$\mathbb{E}\|H\|_{F}^{2}\leq M^{2}(v_{n}+\mu_{n}^{2}),\qquad\mathbb{E}\|H\|_{2}^{2}\leq M^{2}(v_{n}+\mu_{n}^{2}),\qquad\mathbb{E}\|H\|_{2}\leq M\sqrt{v_{n}+\mu_{n}^{2}}.$
Lemma 3.7 converts the entrywise statistics into matrix-scale magnitude bounds. While these bounds do not analytically constrain the spectrum (which we examine numerically in Section 4.3), they rigorously establish the vanishing scale of the operator. Applying this tool to the scalings established in Theorem 3.3 yields the following two corollaries. Corollary 3.8 bounds the typical Hessian magnitude, while Corollary 3.9 incorporates finite-shot noise to quantify the sampling cost.
@@ 3.3 From entrywise variance to matrix-scale bounds and sample complexity
Under the conditions of Theorem 3.3. Suppose further that there exists a constant $c>0$, independent of $n$, such that $|\mathbb{E}_{\rho}[H_{jk}]|^{2}\leq c\operatorname{Var}_{\rho}(H_{jk})$ for all $j,k$. Let $v_{n}$ denote the entrywise variance bound from Theorem 3.3. Then
@@
While Corollary 3.8 bounds the intrinsic curvature scale, the practical bottleneck in VQAs is the signal-to-noise ratio under finite sampling. If the entrywise fluctuations $\sqrt{\operatorname{Var}_{\rho}(H_{jk})}$ are small, resolving them against shot noise becomes computationally prohibitive. The following result quantifies this sampling cost.
Corollary 3.9 (Sample complexity).
Let $\widehat{H}_{jk}(\bm{\theta})$ be the estimator computed via Algorithm 1 using $N$ independent shots per circuit evaluation. Since the parameter-shift rule is a fixed linear combination, the estimator variance satisfies
@@
where $\sigma^{2}\leq 1$ is the single-shot variance bound (given $\|O\|\leq 1$) and $c_{\mathcal{S}}=\sum_{s\in\mathcal{S}_{jk}}a_{s}^{2}$ is a constant determined by the shift rule. Consequently, ensuring a root-mean-square estimation error of $\varepsilon$ requires a shot complexity $N\gtrsim c_{\mathcal{S}}\sigma^{2}/\varepsilon^{2}$.
Theorem 3.3 and Corollary 3.9 reveal a sharp dichotomy in the computational cost of resolving Hessian entries. Targeting an $O(1)$ signal-to-noise ratio requires the shot complexity $N$ to scale inversely with variance. For global objectives, exponential variance suppression dictates $N\gtrsim\exp(2\tilde{\alpha}(r,L)n)$, rendering optimization intractable. In contrast, for termwise local objectives, polynomial decay ensures a feasible scaling $N\gtrsim n/V_{G}(k+2rL)$, preserving statistical accessibility for sub-extensive lightcones.
Algorithm 1 summarizes the finite-shift procedure underlying this analysis, which serves as the basis for the numerical experiments presented in § 4.
Algorithm 1 Finite-complexity estimation of a Hessian entry via parameter-shift
1:Ansatz $U(\bm{\theta})$, objective $C(\bm{\theta})$, indices $(j,k)$, shots NN
2:Unbiased estimate $\widehat{H}_{jk}$ of the Hessian entry $H_{jk}$
[…]
4:for all $s\in\mathcal{S}_{jk}$ do $\triangleright$ Loop over shift set $\mathcal{S}_{jk}$ with coeff. $a_{s}$ (cf. § 2.3)
5:  Prepare $|\psi(\bm{\theta}\oplus s)\rangle$ and measure observable $O$ with NN shots
7:  $\widehat{H}_{jk}\leftarrow\widehat{H}_{jk}+a_{s}\,\widehat{C}(\bm{\theta}\oplus s)$ $\triangleright$ Linear combination preserves unbiasedness
8:end for
9:return $\widehat{H}_{jk}$
@@ 4 Numerical Experiments
This section corroborates the variance-scaling predictions of § 3 and examines their practical implications under finite evaluation complexity. After describing the simulation setting and the estimators used for gradient and Hessian statistics, we report: (i) system-size and depth scaling of representative Hessian-entry variances; (ii) complementary spectral diagnostics of $H(\bm{\theta})$ at initialization; (iii) budget-limited optimization experiments with matched oracle calls; and (iv) an additional VQE benchmark for a $k$-local Hamiltonian, together with quantitative spectral summary metrics linking locality to resolvable curvature.
@@ 4.1 Experimental setup
All simulations were performed in Python using PennyLane [bergholm2018pennylane] with an exact statevector backend, so the reported statistics isolate intrinsic landscape fluctuations. We instantiate (2) as a layered hardware-efficient ansatz: in each layer $\ell$, $V_{\ell}$ is a fixed nearest-neighbor brickwork CNOT pattern and the parametrized gates are single-qubit $R_{y}$ rotations,
@@
For each sampled initialization $\bm{\theta}$ we evaluate gradients and Hessian entries via the parameter-shift identities (7)–(9) on the statevector backend. Unless otherwise stated, the scaling experiments use $n\in\{2,\dots,16\}$, $L\in\{1,\dots,12\}$, and $N_{s}=200$ i.i.d. initializations per configuration $(n,L)$; variances are computed as empirical second central moments over this ensemble.
We estimate variances by the sample variance $\widehat{\operatorname{Var}}$ computed over $N_{s}$ i.i.d. random initializations. Exponential scalings in $n$ are obtained by least-squares fits of $\log\widehat{\operatorname{Var}}$ against $n$ (i.e., $\log\widehat{\operatorname{Var}}(n)\approx a-\alpha n$), and polynomial scalings by least-squares fits of $\log\widehat{\operatorname{Var}}$ against $\log n$ (i.e., $\log\widehat{\operatorname{Var}}(n)\approx b-p\log n$). 95%-percentile confidence intervals for $\widehat{\operatorname{Var}}$ are computed by a nonparametric bootstrap with 500 replicates.
@@ 4.2.1 Dependence on system size
Figure 2 then reports empirical variances of a diagonal and an off-diagonal entry, $\operatorname{Var}_{\rho}(H_{jj})$ and $\operatorname{Var}_{\rho}(H_{jk})$, as functions of $n$, for both the global and local objectives. Unless otherwise stated, results are shown for $(j,k)=(1,2)$; other index choices exhibit the same scaling up to constant prefactors.
For $C_{\mathrm{global}}$ (red markers), both diagonal and off-diagonal entries exhibit clear exponential decay with $n$. Least-squares fits of $\log\operatorname{Var}_{\rho}(H_{jj})$ and $\log\operatorname{Var}_{\rho}(H_{jk})$ versus $n$ over $n$ yield
@@
consistent with Theorem 3.3 under Assumption 3.2. Via Corollary 3.9, this implies that the target resolution required to track typical entrywise curvature becomes exponentially small as $n$ grows.
@@
Figure 2: Hessian-entry variance vs. system size. $\operatorname{Var}[H_{jj}]$ (left) and $\operatorname{Var}[H_{jk}]$ (right) versus $n$ for $C_{\mathrm{global}}$ (red) and $C_{\mathrm{local}}$ (blue); dashed lines show exponential/power-law fits. Error bars: $95\%$ CIs over $N_{s}=200$ initializations.

### v2

@@ 3.2 Main results
Remark 3.2.
Assumption 3.2 is a second-moment concentration condition on the objective value at random initialization. Its role is to serve as an input for the global regime; Lemma 3.3 shows that such concentration is uniform over the finite shift sets used in parameter-shift rules and is preserved under any constant-size finite-shift linear combination.
The assumption holds, for example, when the initialized state ensemble forms an (approximate) projective $2$-design, in which case $\operatorname{Var}_{\rho}[C(\bm{\theta})]$ decays as $\mathcal{O}(2^{-n})$ for $\|O\|\leq 1$ (see Appendix B.2 and Proposition B.3). In architectures where design guarantees are unavailable at the considered depth, (13) can be assessed directly by Monte Carlo estimation of $\operatorname{Var}_{\rho}[C(\bm{\theta})]$; by shift-invariance of $\rho$, the same estimate applies to all shifted evaluations appearing in the parameter-shift rules.
The next lemma formalizes this transference: second-moment concentration of $C(\bm{\theta})$ is uniform over shifts and stable under finite-shift linear combinations.
@@
Let $\rho$ be shift-invariant on the torus, i.e., $\bm{\theta}\sim\rho\Rightarrow\bm{\theta}\oplus s\sim\rho$ for all shifts $s\in\mathbb{R}^{M}$. Fix a finite shift set $\mathcal{S}\subset\mathbb{R}^{M}$ and coefficients $\{w_{s}\}_{s\in\mathcal{S}}\subset\mathbb{R}$, and define $X_{s}:=C(\bm{\theta}\oplus s)$. If $\operatorname{Var}_{\rho}[C(\bm{\theta})]\leq v_{n}$, then the same bound holds uniformly over shifts,
$\operatorname{Var}_{\rho}[X_{s}]=\operatorname{Var}_{\rho}[C(\bm{\theta}\oplus s)]=\operatorname{Var}_{\rho}[C(\bm{\theta})]\leq v_{n}\qquad\text{for all }s\in\mathcal{S},$
@@
$\operatorname{Var}_{\rho}\Big(\sum_{s\in\mathcal{S}}w_{s}\,C(\bm{\theta}\oplus s)\Big)\leq\Big(\sum_{s\in\mathcal{S}}|w_{s}|\Big)^{2}v_{n}.$
Proof 3.4.
Shift-invariance implies $\bm{\theta}$ and $\bm{\theta}\oplus s$ have the same distribution under $\rho$, hence $\operatorname{Var}_{\rho}[C(\bm{\theta}\oplus s)]=\operatorname{Var}_{\rho}[C(\bm{\theta})]$ for all $s\in\mathcal{S}$. For the linear combination, by bilinearity,
$\operatorname{Var}\Big(\sum_{s\in\mathcal{S}}w_{s}X_{s}\Big)=\sum_{s,s^{\prime}\in\mathcal{S}}w_{s}w_{s^{\prime}}\operatorname{Cov}(X_{s},X_{s^{\prime}}).$
@@
$\operatorname{Var}\Big(\sum_{s\in\mathcal{S}}w_{s}X_{s}\Big)\leq v_{n}\sum_{s,s^{\prime}\in\mathcal{S}}|w_{s}||w_{s^{\prime}}|=\Big(\sum_{s\in\mathcal{S}}|w_{s}|\Big)^{2}v_{n}.$
@@
(14) $V_{G}(m):=\max_{v\in V(G)}\big|B_{G}(v,m)\big|.$
For bounded-degree graphs, $V_{G}(m)$ is finite for each fixed $m$; for $D$-dimensional lattices one has $V_{G}(m)=\mathcal{O}(m^{D})$. We are then ready to state our main theorem.
@@
Proof 3.6 (Proof sketch).
@@
For global objectives, Lemma 3.3 transfers the assumed second-moment concentration of $C(\bm{\theta})$ uniformly over the finite shifts; since the parameter-shift rule uses only $O(1)$ evaluations with $O(1)$ coefficients, the exponential rate in $n$ is preserved up to a constant factor, yielding (15).
For local objectives, bounded lightcone growth implies that sufficiently separated local contributions depend on disjoint subsets of parameters; under the product initialization measure $\rho$, their covariances vanish. A dependency-graph argument then gives $\operatorname{Var}_{\rho}(C(\bm{\theta}\oplus s))=\mathcal{O}(V_{G}(k+2rL)/n)$ uniformly in $s$. Plugging this bound into the covariance representation yields (16). Details are deferred to Appendix A.3.
@@
The estimate (16) is an upper bound obtained via a generic dependency-graph argument and is not expected to be exponent-tight in typical instances. The main message of Theorem 3.5 is that the measurement cost of resolving Hessian entries is controlled by lightcone growth: exponential decay can occur for global objectives, while bounded-depth locality yields polynomial bounds.
3.3 From entry-wise variance to matrix norms and sample complexity
While the entry-wise bounds above quantify the typical fluctuation scale, the practical utility of curvature information depends on the number of shots required to separate these fluctuations from measurement noise, and on how they aggregate into matrix-level quantities. Throughout this subsection, $\mathbb{E}_{\rho}[\cdot]$ and $\operatorname{Var}_{\rho}(\cdot)$ denote expectation and variance with respect to initialization $\bm{\theta}\sim\rho$. When finite-shot estimation is involved, we write $\mathbb{E}_{\mathrm{sh}}[\cdot\mid\bm{\theta}]$ and $\operatorname{Var}_{\mathrm{sh}}(\cdot\mid\bm{\theta})$ for expectation and variance over measurement noise (shots), conditional on a fixed parameter value $\bm{\theta}$.
We first record a standard implication of entry-wise second-moment control for matrix norms; a proof is included in Appendix B.3 for completeness.
Lemma 3.9 (Entry-wise second moments imply Frobenius and spectral norm bounds).
$\mathbb{E}\|H\|_{F}^{2}=\sum_{j,k=1}^{M}\mathbb{E}[H_{jk}^{2}],\qquad\mathbb{E}\|H\|_{2}^{2}\leq\mathbb{E}\|H\|_{F}^{2},\qquad\mathbb{E}\|H\|_{2}\leq\sqrt{\mathbb{E}\|H\|_{F}^{2}}.$
In particular, if $\operatorname{Var}(H_{jk})\leq v_{n}$ and $\big|\mathbb{E}[H_{jk}]\big|\leq\mu_{n}$ for all $j,k$, then
$\mathbb{E}\|H\|_{F}^{2}\leq M^{2}(v_{n}+\mu_{n}^{2}),\qquad\mathbb{E}\|H\|_{2}\leq M\sqrt{v_{n}+\mu_{n}^{2}}.$
Lemma 3.9 converts entry-wise second-moment bounds into a global norm scale for the Hessian at initialization. These estimates quantify typical magnitude (e.g., for Hessian–vector products) but do not, by themselves, control conditioning or extremal eigenvalues without additional spectral information. Applying this tool to the scalings established in Theorem 3.5 yields the following two corollaries.
@@ 3.3 From entry-wise variance to matrix norms and sample complexity
Under the conditions of Theorem 3.5. Suppose further that there exists a constant $c>0$, independent of $n$, such that $|\mathbb{E}_{\rho}[H_{jk}]|^{2}\leq c\operatorname{Var}_{\rho}(H_{jk})$ for all $j,k$. Let $v_{n}$ denote the entry-wise variance bound from Theorem 3.5. Then
@@
While Corollary 3.10 bounds the intrinsic curvature scale, the key practical question is the resolution cost in Definition 3.1 under finite-shot sampling.
Corollary 3.11 (Absolute-accuracy shot cost at fixed parameters).
Let $\widehat{H}_{jk}(\bm{\theta})$ be the estimator computed via Algorithm 1 using $N$ independent shots per circuit evaluation. Since the parameter-shift rule is a fixed linear combination, the shot-noise contribution to the estimator variance satisfies
@@
where $\sigma^{2}\leq 1$ when $\|O\|\leq 1$ and $c_{\mathcal{S}}:=\sum_{s\in\mathcal{S}_{jk}}w_{s}^{2}$ depends only on the chosen shift rule. Consequently, achieving $\mathrm{RMSE}\leq\varepsilon$ for estimating $H_{jk}(\bm{\theta})$ at fixed $\bm{\theta}$ requires $N\gtrsim c_{\mathcal{S}}\sigma^{2}/\varepsilon^{2}$. This bound concerns absolute accuracy at fixed $\bm{\theta}$; resolvability relative to the initialization scale is addressed in Definition 3.1 and Corollary 3.12.
Corollary 3.12 (Resolution shot requirement).
Under the assumptions of Corollary 3.11, the resolution criterion in Definition 3.1 is ensured whenever
$N\gtrsim\frac{c_{\mathcal{S}}\sigma^{2}}{\eta\operatorname{Var}_{\rho}(H_{jk})}.$
In particular, combining this with the variance bounds in Theorem 3.5 yields exponential scaling of the resolution cost for global objectives and polynomial scaling for term-wise local objectives (under sub-extensive backward-lightcones).
Combining Theorem 3.5 with Corollary 3.12 shows that the shot requirement for resolving Hessian entries is governed by the initialization variance $\operatorname{Var}_{\rho}(H_{jk})$. For global objectives, exponential decay of $\operatorname{Var}_{\rho}(H_{jk})$ forces an exponentially growing resolution cost in $n$. For term-wise local objectives in bounded-depth circuits, $\operatorname{Var}_{\rho}(H_{jk})$ is bounded by a polynomial controlled by backward-lightcone growth, leading to polynomial resolution cost when the lightcone remains sub-extensive.
@@ 4 Numerical Experiments
This section validates the variance-scaling predictions of § 3 and studies the empirical implications of these scaling laws under finite-shot sampling. After describing the simulation setting and variance estimators, we present: (i) system-size and depth scaling of representative Hessian-entry variances, including a VQE benchmark for a $k$-local Hamiltonian; (ii) spectral diagnostics of $H(\bm{\theta})$ at initialization, including histogram estimates and scalar summary metrics; and (iii) finite-shot optimization trajectories with matched oracle calls to illustrate the practical effect of the resolution criterion (Definition 3.1) near initialization.
@@ 4.1 Experimental setup
All simulations were performed in Python using PennyLane [bergholm2018pennylane] with an exact statevector backend, which ensures that the reported statistics isolate intrinsic landscape fluctuations. We instantiate (2) as a layered hardware-efficient ansatz: in each layer $\ell$, $V_{\ell}$ is a fixed nearest-neighbor brickwork CNOT pattern and the parametrized gates are single-qubit $R_{y}$ rotations,
@@
For each sampled initialization $\bm{\theta}$ we evaluate gradients and Hessian entries via the parameter-shift identities (6)–(8) on the statevector backend. Unless otherwise stated, the scaling experiments use $n\in\{2,\dots,16\}$, $L\in\{1,\dots,12\}$, and $N_{s}=200$ i.i.d. initializations per configuration $(n,L)$; variances are computed as empirical second central moments over this ensemble. Results are shown for $(j,k)=(1,2)$; other index choices exhibit the same scaling up to constant prefactors.
We estimate $\operatorname{Var}_{\rho}(\cdot)$ by the sample variance $\widehat{\operatorname{Var}}$ computed from $N_{s}$ i.i.d. random initializations $\bm{\theta}\sim\rho$. To extract scaling with system size $n$, we fit exponential laws by least squares on the log scale and polynomial laws by least squares in log–log form. When reported, error bars denote 95% bootstrap confidence intervals for $\widehat{\operatorname{Var}}$ obtained by resampling the $N_{s}$ initializations.
@@ 4.2.1 Dependence on system size
Fig. 2 then reports empirical variances of a diagonal and an off-diagonal entry, $\operatorname{Var}_{\rho}(H_{jj})$ and $\operatorname{Var}_{\rho}(H_{jk})$, as functions of $n$, for both the global and local objectives. For $C_{\mathrm{global}}$ (red markers), both diagonal and off-diagonal entries exhibit clear exponential decay with $n$. Least-squares fits of $\log\operatorname{Var}_{\rho}(H_{jj})$ and $\log\operatorname{Var}_{\rho}(H_{jk})$ yield
@@
consistent with Theorem 3.5 under Assumption 3.2. Via Corollary 3.11, this implies that the target resolution required to track typical entry-wise curvature becomes exponentially small as $n$ grows.
@@
Figure 2: Hessian-entry variance vs. system size. $\operatorname{Var}[H_{jj}]$ (left) and $\operatorname{Var}[H_{jk}]$ (right) versus $n$ for $C_{\mathrm{global}}$ (red) and $C_{\mathrm{local}}$ (blue).